RC Update and Run Plan for Aug 9, 2020

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 4:15 PM, remote

Sunday Aug 9 2020.

General updates

  • Had trippy beam at the start of swing. After some tuning, proceeded to begin BCM calibration. Only got through one cycle (too long and invasive to Hall D). [1]
  • Reduced run current to 27 uA to improve SHMS live time. Fairly smooth running after 8pm when 1L13 found to be draining a bit of energy from C beam line. [2]
  • No beam initially due to beam steering issues (BLM trips) in 7S at the start of day shift.
  • Access in progress to NL for 1L20 vacuum issues. [3],[4],[5],[6],[7]

Pending problem

  • HMS dipole throws current lead mass flow interlock 3 times today and daily the past few days. Steve L will do some re-programming of this fault on Tuesday: [8]

Recent polarization measurements

  • 43.9% at 11:30am [9]

Current Run Plan

Normal Production running

  • Prerequisites
    • Make sure the raster size is correct (4-4.5)
    • Monitor beam parameters/configuration very often and post an entry like this at the beginning of your shift and when run conditions have changed
    • Running on cell targets: raster always ON (4-4.5, size will be determined by the scraping checks), orbit locks ON, current ramp rate 1 muA/sec
    • Always take 120Hz of EDTM
    • Fill the BTA with realistic numbers. If the desired beam current is not delivered prorate your ABUs!
    • Change the IHWP every 4 hours when running on glass cells
  • Run plan
    • Re-establish beam [10].
    • Production on 3He; if not able to reach requested current (27 uA), pro-rate the ABUs

Notes for tomorrow's 7:45 meeting

Future

  • Opportunistic target polarization calibration and studies on Tuesday-Wednesday during beam off and Moller measurement (priority is beam)
  • Spin polarization change 270->90 on Tuesday start of RF recovery.
  • Moller measurement on Wednesday
  • Plan for BCM calibration when C has priority. Need two more cycles (disruptive to D due to shared slit).
  • After ~ 10 days of running Hall C will go to 1 pass for ~3 days

Hall Access

  • Planning for Simona's SoLID Cerenkov access on Tuesday, will need to coordinate to rotate HMS to 29 deg (Steve L), crane work, etc. ~4-5 hours work. Walter, Jay, Steve notified. Working with RADCON to be swept first thing Tuesday.
  • Tgt people: swap EPR photodiode on Tuesday?

When there is a down:

Notices in Effect for Shift Crew

  • Beam positions: [11]
  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • Spectrometer GUI (go_magnets) should always be visible on Shift Leader computer, hcdesk3
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
    • If any of the Spectrometers magnets trip, take a screenshot of the PSU window and check the hall for smoke or water with the camera before resetting the interlock and post it to hclog.
  • Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
  • Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
  • Ramp rate lock instructions are here and added to checklist[12]

Planned Accelerator Configuration Changes

  • Bluejeans remote connection information:
 To join the Meeting:
 https://bluejeans.com/861439920
 To join via Room System:
 Video Conferencing System: bjn.vc -or-199.48.152.152
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links